# DoubleTake: Geometry Guided Depth Estimation

### v1

@@ 5 Conclusion
We introduced a system for depth estimation and 3D reconstruction which can take as input, where available, previously-made estimates of the scene’s geometry. Our carefully-designed architecture takes as input geometrical hints, and makes high quality depth estimates even when these hints are not available. We showed how our method can make use of hints from the near-term for instantaneous depths in new environments, and also hints from the past when estimating depths in previously visited locations. We evaluated on a range of challenging datasets where we have obtained state-of-the-art depths and reconstructions.
Figure 9: Our method generalizes to new domains e.g. this outside scene, casually captured with a smartphone. Poses are from ARKit [1].

### v2

@@ 5 Conclusion
We introduced a system for depth estimation and 3D reconstruction which can take as input, where available, previously-made estimates of the scene’s geometry. Our carefully-designed architecture takes as input geometrical hints, and can make high quality depth estimates even when these hints are not available. We demonstrated how our method can make use of hints from the near-term for instantaneous depths in new environments, and also hints from the more distant past, e.g. when estimating depths in previously visited locations. We evaluated on a range of challenging datasets where we have obtained state-of-the-art depths and reconstructions.